### v2

# GOProteinGNN: Leveraging Protein Knowledge Graphs for Protein Representation Learning

Abstract.
Proteins play a vital role in biological processes and are indispensable for living organisms. Accurate representation of proteins is crucial, especially in drug development. Recently, there has been a notable increase in interest in utilizing machine learning and deep learning techniques for unsupervised learning of protein representations. However, these approaches often focus solely on the amino acid sequence of proteins and lack factual knowledge about proteins and their interactions, thus limiting their performance. In this study, we present GOProteinGNN, a novel architecture that enhances protein language models by integrating protein knowledge graph information during the creation of amino acid level representations. Our approach allows for the integration of information at both the individual amino acid level and the entire protein level, enabling a comprehensive and effective learning process through graph-based learning. By doing so, we can capture complex relationships and dependencies between proteins and their functional annotations, resulting in more robust and contextually enriched protein representations. Unlike previous methods, GOProteinGNN uniquely learns the entire protein knowledge graph during training, which allows it to capture broader relational nuances and dependencies beyond mere triplets as done in previous work. We perform a comprehensive evaluation on several downstream tasks demonstrating that GOProteinGNN consistently outperforms previous methods, showcasing its effectiveness and establishing it as a state-of-the-art solution for protein representation learning.
Knowledge Graphs, Protein Language Models, Bioinformatics
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: 31th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 3–7, 2025; Toronto, ON, Canadaisbn: 978-1-4503-XXXX-X/18/06ccs: Applied computing Bioinformaticsccs: Computing methodologies Artificial intelligence
1. Introduction
Proteins play a vital role in biological processes and are central to understanding the mechanisms of living organisms. In the pursuit of unraveling their complexity, the challenge of accurately representing proteins has gained considerable attention. Accurate representation holds paramount significance, especially for drug development (Liu et al., 2024).
Amino acids serve as the fundamental building blocks of proteins. In recent years, there has been a surge of interest in employing deep learning techniques to create unsupervised amino acid representations that can be used to represent a protein (Rives et al., 2019; Rao et al., 2021). Prevailing approaches have predominantly focused on unraveling amino acid relationships using text-based methodologies, such as ProtBert (Elnaggar et al., 2022). The amino acid representations are then averaged to create the protein representation.
Recently, several approaches have attempted to directly create a protein representation by integrating external information about proteins, rather than relying solely on the amino acid representations. Those methods, such as KeAP (Zhou et al., 2023), harness vast protein knowledge graphs. These graphs contain diverse information about interactions the protein participated in, the processes in which it is involved, etc. The factual knowledge is conveyed through text descriptions. Figure 3 illustrates an example of such a protein knowledge graph, showcasing the intricate relations between proteins and Gene Ontology (GO) terms. However, current state-of-the-art (SOTA) methods still encounter challenges in effectively incorporating this extensive information.
Existing paradigms, at times, reduce the intricate protein knowledge graph to mere triplets (Protein, Relation, GO Term), neglecting the complex tapestry of the knowledge graph (Zhou et al., 2023). These methods have overlooked the importance of capturing a holistic graph representation. For instance, although they may encode annotations, crucial information such as broader relational nuances and contextual dependencies within the graph may be lost unless the entire graph structure is considered. The aforementioned models also incorporate knowledge post-encoder, either via loss objectives or pre-training enrichment (Zhou et al., 2023; Elnaggar et al., 2022). Language models often struggle to accurately embed knowledge, particularly long-tail knowledge (Kandpal et al., 2022). This limitation hinders their ability to effectively incorporate external knowledge into protein representations during the encoder stage. Moreover, these models have predominantly concentrated on either amino acid level learning (Zhou et al., 2023; Elnaggar et al., 2022) or protein level learning (Zhang et al., 2022; Lam et al., 2023), with none addressing both simultaneously.
In this work, we present GOProteinGNN, a novel architecture that integrates amino acid level and protein level representations within a single framework. Our method addresses existing limitations by comprehensively utilizing the entire protein knowledge graph structure during pre-training, ensuring the capture of intricate relationships. The main technical advancement of GOProteinGNN lies in its ability to capture broader relational nuances and dependencies between protein attributes, which existing models often overlook. In a specific context, when an attribute is connected to another attribute, determining how to embed these connections during protein representation learning is a complex challenge. Compared to GOProteinGNN, which addresses this limitation, KeAP (Zhou et al., 2023) and previous studies have failed to account for a holistic view of the protein’s interactions within a biological context, potentially leading to a fragmented understanding of its functional significance.
A key innovation of our approach is the integration of knowledge graph information into protein language models (PLMs), using our novel Graph Neural Networks (GNN) Knowledge Injection (GKI) mechanism, which leverages the [CLS] token representation during the encoder stage. This unique integration enables holistic learning of protein representations, as reflected in our SOTA results over extensive real-world bioinformatics tasks.
Specifically, GOProteinGNN interprets amino acid sequences as coherent “sentences” that describe proteins, learning representations for individual amino acids in its initial encoder layers. The GKI layer then assigns representations to protein nodes in the protein knowledge graph and incorporates knowledge into these representations by leveraging neighboring representations through a Relational Graph Convolution Network (RGCN) (Schlichtkrull et al., 2017). These refined representations are processed by subsequent encoder layers, generating comprehensive amino acid representation. This process alternates between amino acid level representations and graph representation, enhancing the overall representation by allowing new insights to spread across the amino acid sequence and ensuring a thorough encoding that incorporates both sequential and graph-enhanced information.
We perform an empirical evaluation over numerous protein tasks, spanning a broad spectrum of biological domains, comparing to superior methods and achieving SOTA performance across various benchmarks. We study the performance of our method through ablation tests.
The contributions of this work are threefold: (1) We present a novel algorithm for protein representation learning that allows the integration of knowledge graph information into PLMs during the encoder stage. This enables a comprehensive and effective learning process through graph-based learning. (2) We introduce a novel GNN Knowledge Injection (GKI) mechanism that enables protein knowledge information to propagate through the amino acid sequence representations. We leverage the [CLS] token representation to integrate the GKI with the PLM in a novel manner. (3) We present an extensive empirical evaluation of our work over numerous real-world bioinformatics tasks establishing SOTA results. Also, we contribute our code to the community  [^1] https://github.com/kalifadan/GOProteinGNN.
Figure 1. An example of a Protein Knowledge Graph. The upper figure illustrates a protein with associated biological knowledge, while the lower figure depicts the corresponding knowledge graph. The central node, Q14028, represents the protein connected to multiple GO terms through matching relations (edges), providing insights into Molecular Function, Cellular Component, and Biological Process [^3] https://www.zjukg.org/project/ProteinKG65.
(a) The GKI layer.
(b) The GOProteinGNN architecture.
Figure 2. The GOProteinGNN pre-training architecture. The model incorporates a protein’s knowledge graph with relations and GO terms. It uses encoder layers to process the amino acid sequence, producing amino acid representations and a protein representation (which is represented by the [CLS] representation). Then, the GKI refines the protein’s representation with knowledge graph information using graph learning techniques, yielding a knowledge-enhanced protein representation that captures essential biological contexts and interactions. Subsequent encoder layers further enhance the amino acid representation, while the model performs masked language modeling (MLM) to restore masked amino acids.
2. Related Work
2.1. Proteins Representation Learning
Proteins are essential for living organisms and have vital roles in biological processes. Accurate representation of proteins is crucial, particularly in fields such as drug development. In recent years, there has been a surge of interest in applying machine learning and deep learning techniques to learn protein representations in an unsupervised manner. The popularity gain of Transformers (Vaswani et al., 2017) introduced many novel PLMs such as ESM-1b (Rives et al., 2019), ESM-2 (Lin et al., 2023), MSA Transformer (Rao et al., 2021), ProteinBERT (Brandes et al., 2021), ProteinLM (Xiao et al., 2021) and ProtTrans (Elnaggar et al., 2022). These models were trained on data from UniRef (Suzek et al., 2007) which contains hundreds of billions of amino acids.
A protein’s biological mechanism is determined by its three-dimensional (3D) native structure (Dill and MacCallum, 2012). AlphaFold (Jumper et al., 2021; Senior et al., 2020) is a recent model predicting the 3D structure of a protein. A common SOTA method that resembles the AlphaFold architecture but tuned for protein representation is ESM (Rives et al., 2019; Lin et al., 2023). Unlike AlphaFold, which primarily focuses on 3D structure prediction and relies on multiple sequence alignments and template structures, ESM requires only a single input sequence (i.e., protein) and has demonstrated similar performance across various tasks. Our comparative analysis focuses on ESM, emphasizing its utilization of single-sequence input and acknowledging its performance similarity to AlphaFold.
2.2. Knowledge-Enhanced Language Models
Prior research has demonstrated the effectiveness of leveraging knowledge in language models to enhance their performance (Zhang et al., 2019; Sun et al., 2019). Specifically, KEPLER (Wang et al., 2019) endeavored to integrate factual knowledge through knowledge embedding optimization, while KG-BERT (Yao et al., 2019) and K-BERT (Liu et al., 2019) treated knowledge graph triples as textual sequences or transformed the original input sentence into a knowledge-rich sentence tree, respectively. In contrast, our algorithm differs significantly, as it considers graph topology and diverse node types present in protein knowledge graphs.
While PLMs traditionally learn protein representations from sequences, ProtST (Xu et al., 2023) was trained on biomedical text descriptions. Similarly, OntoProtein (Zhang et al., 2022) integrated external knowledge graphs into protein pre-training, focusing on Gene Ontology (GO) term and protein entity alignment. KeAP (Zhou et al., 2023) explored more granular knowledge graph interaction for improved protein decoding. OntoProtein’s contrastive learning and masked modeling were enriched by KeAP’s focus on the masked language modeling objective, achieving superior performance in various tasks. Otter-Knowledge (Lam et al., 2023) extended knowledge graph integration beyond proteins, incorporating diverse entities into a comprehensive knowledge graph.
The aforementioned models also incorporate knowledge post-encoder, either via loss objectives or pre-training enrichment (Zhou et al., 2023; Elnaggar et al., 2022). Language models often struggle to accurately embed knowledge, particularly long-tail knowledge (Kandpal et al., 2022). This limitation hinders their ability to effectively incorporate external knowledge into protein representations during the encoder stage.
Both OntoProtein (Zhang et al., 2022) and KeAP (Zhou et al., 2023), together with other existing models, have predominantly focused on either amino acid level learning or protein level learning. Also, they reduce the intricate protein knowledge graph to mere triplets, neglecting the intricate tapestry of the knowledge graph. GOProteinGNN sets itself apart by uniquely encompassing both levels of representation within a single framework thus learning the entire protein knowledge graph and allowing it to capture broader relational nuances and dependencies beyond mere triplets.
We specifically target protein-related tasks, capturing critical biological contexts and interactions unique to proteins.
3. Methods
We present the GOProteinGNN architecture (Fig. 2), which extends the capabilities of PLMs by incorporating knowledge graph information into the models, using our proposed GKI layer. We leveraged the [CLS] token representation, to integrate between the proposed GKI and the PLM in a novel manner. This integration of language model-based representation learning with knowledge graph processing allows GOProteinGNN to generate enriched protein representations that capture essential biological contexts and interactions. The detailed algorithm is presented in Supplementary Material A.
3.1. Protein Knowledge Graph
GOProteinGNN is trained on a knowledge graph consisting of two different types of nodes: Proteins and GO-Terms, with Relations that act as edges. Proteins are represented as a sequence of amino acids, while the relation and the GO term are terms of factual knowledge described in natural language. The GO terms cover three aspects of biology: Molecular Function, Cellular Component, and Biological Process.
Protein Knowledge Graph Definition
Let $V_{proteins}$ be a set containing proteins (amino acid sequences). The GKI leverages the accompanying descriptions of GO terms extracted from the protein knowledge graph. These descriptions are processed through a language encoder  [^4] We use PubMedBERT (Gu et al., 2020) to generate comprehensive knowledge representations. To bridge the gap between text and protein, the GKI utilizes affine transformation to project those representations to the same space. These representations in the protein knowledge graph are denoted by $V_{terms}$. We define the knowledge graph $G=(V,E)$ where: $V=V_{proteins}\cup V_{terms}$ and $E=\{(u,v)\mid u\in V_{proteins},v\in V_{terms},(u,v)\in Relations\}$.
3.2. GKI Layer
We introduce a method called GNN Knowledge Injection (GKI), inspired by GNN (Scarselli et al., 2009), a powerful neural network designed for graph-structured data (Liu et al., 2021; Agiollo and Omicini, 2022). GKI is particularly suitable for effectively learning from protein knowledge graphs, analogous to applications in other domains (Sun, 2022; Wang et al., 2021). By utilizing a protein knowledge graph, denoted as $G$, the GKI layer generates an updated protein representation that incorporates information from the knowledge graph into the input representation.
GNN Method
In response to the complex diversity of edge types in the protein knowledge graph, we select a GNN that accommodates the multifaceted edge information. Recognizing that conventional GCNs (Kipf and Welling, 2017) lack the capacity to fully exploit this characteristic, we turn to the Relational Graph Convolution Network (RGCN) (Schlichtkrull et al., 2017), a specialized variant of GNN. By doing so, we can effectively harness the rich information present in the knowledge graph to refine and enhance the protein representation. This approach allows us to leverage the graph structure of the protein knowledge graph and to capture complex relationships and dependencies between proteins and their variable number of functional annotations, leading to more powerful and contextually enriched protein representations, a capability not attainable through other methods, such as representation concatenation. Moreover, other graph algorithms could also be applied. We compare different graph algorithms in Section 6.1.
3.2.1. GKI Procedure
Given a protein representation (i.e., the [CLS] representation) $z_{protein}\in\mathbb{R}^{D}$ corresponding to node $i$ in the knowledge graph $G$, set: $h_{i}^{(0)}=z_{protein}$.
The given protein is connected to multiple nodes in the knowledge graph $G$, such as the GO terms nodes ($V_{terms}$). The $(l+1)$ RGCN layer performs the following update to the node representation $z_{protein}=h_{i}^{(l)}$:
$h_{i}^{(l+1)}=\sigma\left(\sum_{r\in R}\sum_{j\in N_{r}(i)}\frac{1}{c_{i,r}}W_{r}^{(l)}h_{j}^{(l)}+W_{0}^{(l)}h_{i}^{(l)}\right)$
where $h_{i}^{(l+1)}\in\mathbb{R}^{D}$ is the new node representation, $\sigma(\cdot)$ is the activation function, $N_{r}(i)$ is the set of neighbor nodes of node $i$ connected by relation type $r$, $R$ is the set of all relation types in the graph, $h_{j}^{(l)}\in\mathbb{R}^{D}$ is the node representation of the neighbor node $j$ at layer $l$, $c_{i,r}$ is a normalization term for relation $r$ specific to node $i$, and $W_{r}^{(l)}\in\mathbb{R}^{D\times D}$ and $W_{0}^{(l)}\in\mathbb{R}^{D\times D}$ are learnable weight matrices associated with relation $r$ and the self-loop, respectively.
The new protein representation after the $L$ RGCN layers is given by: $z^{{}^{\prime}}_{protein}=h_{i}^{(L)}$ where $z^{{}^{\prime}}_{protein}\in\mathbb{R}^{D}$ represents the updated protein representation that captures essential biological contexts and interactions.
3.3. GOProteinGNN Algorithm
3.3.1. Training Phase
Following the approach of KeAP (Zhou et al., 2023), we use a large pre-trained PLM (ProtBert (Elnaggar et al., 2022)) as our initialization for the encoder layers’ weights, leveraging its contextual understanding of proteins. The usage in ProtBert is a common practice (Zhang et al., 2022; Zhou et al., 2023) and allows us to show the significance of the injection method compared to the baselines in a fair manner. Alternatively, other advanced encoders can be used in the training algorithm.
During the training stage, each amino acid sequence (i.e., protein) is passed through multiple encoder layers, yielding an intermediate representation of the amino acid sequence, denoted by $z_{p}\in\mathbb{R}^{L_{p}\times D}$, where $L_{p}$ is the sequence length and $D$ is the feature dimension (line 5 in the algorithm).
In language models, including a specialized classification token ([CLS]) at the beginning of the sequence is a common practice. The final hidden state of this token serves as a comprehensive sequence representation for various classification tasks (Devlin et al., 2019). In our framework, we uniquely leverage the [CLS] token representation to integrate external knowledge into our representation framework and enhance the protein representation.
Within the context of a given encoder layer denoted by $l$, a GKI layer takes the input representation $[CLS]_{l}$ along with the protein knowledge graph denoted by $G$, resulting in a refined representation $[CLS]^{\prime}_{l}$ (line 9):
$[CLS]^{\prime}_{l}=GKI([CLS]_{l},G)\in\mathbb{R}^{D}$
Where the GKI utilizes the entirety of the knowledge graph during each injection process.
Subsequently, this refined [CLS] token representation assumes the role of the new [CLS] token in the ensuing encoder layer. Then, using the self-attention (Vaswani et al., 2017) of the encoder layer, the refined representation propagates and refines the representations of all other amino acids within the sequence, resulting in a new representation of the amino acid sequence (line 7).
This mechanism enables GOProteinGNN to achieve a holistic learning process, harmonizing amino acid level representations with protein level representations. The outcome is an integration of contextually enriched representations that intricately capture both sequential insights and the graph-enhanced knowledge pertaining to proteins.
As the protein representation undergoes transformations across the encoder layers, our aim is to uniformly integrate knowledge a total of $k$ times across the $L$ encoder layers. Therefore, we add a GKI layer after every $L\div k$ encoder layer (line 8). Unlike prior language models that utilize the [CLS] token only in the final output (Devlin et al., 2019), our architecture iteratively integrates knowledge across the encoder layers, ensuring a more consistent and dynamic integration of knowledge throughout the representation process.
Method Short Range Medium Range Long Range
P@L P@L/2 P@L/5 P@L P@L/2 P@L/5 P@L P@L/2 P@L/5
LSTM 0.26 0.36 0.49 0.20 0.26 0.34 0.20 0.23 0.27
ResNet 0.25 0.34 0.46 0.28 0.25 0.35 0.10 0.13 0.17
Transformer 0.28 0.35 0.46 0.19 0.25 0.33 0.17 0.20 0.24
ProtBert (Elnaggar et al., 2022) 0.30 0.40 0.52 0.27 0.35 0.47 0.20 0.26 0.34
ESM-1b (Rives et al., 2019) 0.38 0.48 0.62 0.33 0.43 0.56 0.26 0.34 0.45
ESM-2 (Lin et al., 2023) 0.40 0.50 0.62 0.35 0.44 0.56 0.27 0.35 0.45
KeAP (Zhou et al., 2023) 0.41 0.51 0.63 0.36 0.45 0.54 0.28 0.35 0.43
GOProteinGNN $\textbf{0.44}^{*}$ $\textbf{0.54}^{*}$ $\textbf{0.66}^{*}$ $\textbf{0.39}^{*}$ $\textbf{0.50}^{*}$ $\textbf{0.60}^{*}$ $\textbf{0.30}^{*}$ $\textbf{0.39}^{*}$ $\textbf{0.48}^{*}$
Table 1. Results on amino acid contact prediction task. Short-range, medium-range, and long-range contacts are contacts between positions that are separated by 6 to 11, 12 to 23, and 24 or more positions, respectively. P@L, P@L/2, and P@L/5 denote the precision scores calculated upon top L (i.e., L most likely contacts), top L/2, and top L/5 predictions, respectively. Statistically significant results with $p<0.05$ using a t-test are marked with an asterisk (*). The best result is highlighted in bold.
3.3.2. Masked Language Modeling Training Objective
For each input protein, 20% of the amino acids are randomly masked. Each masked amino acid $a_{(j)}$ has an 80% chance of being masked for prediction, 10% chance of replacement with a random amino acid, and 10% chance of remaining unchanged. Suppose the number of masked amino acids is $M$, the training objective $\mathcal{L}_{MLM}$ is to minimize:
$\mathcal{L}_{MLM}=-\sum_{j=0}^{M-1}\log P(a_{(j)}|z_{p},\Theta)$
where $z_{p}$ is the protein representation and $\Theta$ are the parameters of the GOProteinGNN model.
3.3.3. Inference Phase
During inference, as factual knowledge is not available, and to eliminate its dependency, we introduce a weight-less injection mechanism. Specifically, GOProteinGNN omits the GKI components utilized solely for pre-training. Furthermore, task-specific classification heads are incorporated to enable predictions for these specific tasks. This separation between pre-training and inference allows GOProteinGNN to generalize effectively across various protein tasks, including those involving unseen proteins.
4. Empirical Evaluation
In this section, we outline our empirical setting. We performed extensive analysis on both intrinsic (e.g., semantic similarity inference of proteins) and extrinsic tasks (e.g., protein-protein interactions), demonstrating the value of the novel representation we provide in real-world bioinformatics tasks. Detailed experimental settings, such as implementation details and computational complexity, are provided in the Supplementary Material C. For pre-training, we employed the ProteinKG25 dataset (Zhang et al., 2022), consistent with all knowledge-enhanced baseline methods (See Supplementary Material B for full details). Furthermore, the limitations of knowledge graphs and potential mitigation strategies are discussed in Supplementary Material B.0.1.
4.1. Tasks
We present an extensive empirical evaluation of our method across a wide range of protein-level and amino acid-level representation tasks. We use Tasks Assessing Protein Embeddings (TAPE) (Rao et al., 2019) as a common benchmark to evaluate protein representation learning. TAPE is a set of biologically relevant semi-supervised learning tasks, such as Contact Prediction and Remote Homology Detection, spread across different domains of protein biology. Furthermore, we tested our representations on the Semantic Similarity Inference task from PROBE (Unsal et al., 2022) and on Protein-Protein Interactions (PPI) from GNN-PPI (Lv et al., 2021). All tasks were divided into specific training, validation, and test splits to ensure that each task tested a biologically relevant generalization that transferred to real-life scenarios. See Supplementary Material D for comprehensive task details and Supplementary Material E for an analysis of additional tasks.
4.2. Baselines
Incorporating SOTA models as our baselines, we used ProtBert (Elnaggar et al., 2022) and ESM-1b (Rives et al., 2019) as benchmarks, following KeAP (Zhou et al., 2023) for fair comparison. Furthermore, we added the SOTA ESM-2 (Lin et al., 2023), which improves the ESM-1b model. These models are considered the top-performing models in the absence of factual knowledge integration. Furthermore, we subjected our analysis to the scrutiny of the most recent and advanced knowledge-driven pre-training model, KeAP (Zhou et al., 2023). We exclude the results of OntoProtein (Zhang et al., 2022) in this paper, as it was presented by KeAP (Zhou et al., 2023) that it is significantly inferior on a dataset without leakage (See Supplementary Material B). For each specific task, we followed the convention of selecting SOTA methods, just as prior research has done (See Supplementary Material C.1 for full details on task-specific baselines).
5. Empirical Results
In this section, we present our empirical results. We validate the statistical significance of performance differences, using a one-tailed paired t-test with a 95% confidence level ($p<0.05$), comparing observations from the tested models. Significant results are marked with an asterisk (*) in the tables.
5.1. Contact Prediction
In Table 1, we present GOProteinGNN’s performance on the contact prediction task compared to six baseline methods. GOProteinGNN significantly outperformed all baseline methods across all categories and metrics. In particular, GOProteinGNN showed superior performance compared to ProtBert (Elnaggar et al., 2022) and ESM-2 (Lin et al., 2023), which lack knowledge graph information. This highlights the importance of incorporating knowledge during representation learning, especially for pairwise amino acid tasks (e.g., contact prediction task), where success depends on understanding the proteins’ functions. Moreover, KeAP (Zhou et al., 2023) incorporates a knowledge post-encoder, limiting its capacity to seamlessly integrate external knowledge into protein representations, as reflected in the performance gains of GOProteinGNN over KeAP. These findings highlight the significant impact of incorporating knowledge graph information in GOProteinGNN via the GKI layers, leading to superior performance in the contact prediction task.
5.2. Semantic Similarity Inference
Table 2 presents the results of the semantic similarity inference task. GOProteinGNN outperformed all baselines significantly in capturing semantic similarities, achieving an increase of 23.8% compared to the second-best model. Biological processes play a crucial role in many drug development processes, making this category particularly important as it pertains directly to drug development. The success in this task underscores GOProteinGNN’s efficacy in acquiring enriched representations that encompass the complex biological processes in which proteins are involved, due to the whole protein knowledge graph learning, surpassing the capabilities of alternative approaches. Furthermore, KeAP (Zhou et al., 2023) showed inferior performance compared to ESM-1b (Rives et al., 2019), which does not use factual knowledge. This result is surprising given that both GOProteinGNN and KeAP utilize the same factual knowledge repository, yet KeAP’s method of knowledge injection failed to outperform the model that does not use knowledge at all.
Method Biological Process
MSA Transformer 0.31
ProtT5-XL 0.21
ProtBert 0.35
ESM-1b 0.42
ESM-2 0.41
KeAP 0.41
GOProteinGNN $\textbf{0.52}^{*}$
Table 2. Results on semantic similarity inference task. The Spearman’s rank is the evaluation metric. Statistically significant results with $p<0.05$ using a t-test are marked with an asterisk (*). The best result is highlighted in bold.
5.3. Protein-Protein Interaction Identification
Table 3 presents the results of the protein-protein interaction (PPI) identification task. GOProteinGNN showed strong performance across the board, achieving the highest F1 scores across all datasets with statistical significance, which can be beneficial for drug discovery or personalized medicine. In particular, GOProteinGNN obtained SOTA results with an increase of approximately 2.1% on the SHS27K dataset, and 2.7% on the SHS148K dataset compared to KeAP (Zhou et al., 2023), with statistical significance.
In conclusion, the results highlight the strengths of our GOProteinGNN, which enables a better biological understanding of the protein, leading to success in predicting interactions between proteins.
Method SHS27K SHS148K
DNN-PPI 48.09 57.40
DPPI 41.43 52.12
PIPR 44.48 61.83
GNN-PPI 63.81 71.37
ProtBert 70.94 70.32
ESM-1b 74.92 77.49
ESM-2 75.05 77.19
KeAP 78.58 77.22
GOProteinGNN $\textbf{80.24}^{*}$ $\textbf{79.27}^{*}$
Table 3. Results on protein-protein interaction (PPI) identification task. The table displays the results in two datasets: SHS27K and SHS148K. The breadth-first search (BFS) method is used. The F1 score is the evaluation metric. Statistically significant results with $p<0.05$ using a t-test are marked with an asterisk (*). The best result is highlighted in bold.
5.4. Remote Homology Detection
In Table 4, we present GOProteinGNN’s performance compared to seven baseline methods on the remote homology detection task. Despite the challenging nature of the task, due to the large number of potential protein folds, GOProteinGNN significantly outperformed all methods, demonstrating its effectiveness in handling the complexities of remote homology detection. Specifically, GOProteinGNN achieved an advantage over KeAP (Zhou et al., 2023), further affirming its ability to capture essential biological contexts and interactions which is crucial for predicting the protein fold structure, as it processed the whole protein knowledge graph. As remote homology detection is a sequence-level classification task, we can see that both KeAP (Zhou et al., 2023) and ProtBert (Elnaggar et al., 2022) achieved equal results, suggesting KeAP fails to leverage the factual knowledge in the protein level in this task. This task highlights the novelty of our algorithm which allows for the integration of information at both the individual amino acid level and the entire protein level.
Method Homology
LSTM 0.26
ResNet 0.17
Transformer 0.21
ProtBert 0.29
ProteinBert 0.22
ESM-1b 0.11
ESM-2 0.13
KeAP 0.29
GOProteinGNN $\textbf{0.31}^{*}$
Table 4. Results on remote homology detection task. The evaluation metric is accuracy. Statistically significant results with $p<0.05$ using a t-test are marked with an asterisk (*). The best result is highlighted in bold.
6. Ablation Tests
We conducted ablation tests on all tasks to comprehensively assess the performance of our proposed algorithm. Consistent trends were observed across tasks; hence, we focus on the results for contact prediction and PPI tasks, which are widely studied benchmarks.
6.1. GKI Ablation
The ablation test for the GKI component is presented in Table 5. The full model results highlight the effectiveness of GOProteinGNN’s model components, where the model demonstrated superior performance across all metrics, underscoring its pivotal role. Ignoring relations through the GKI component implies using GCN (Kipf and Welling, 2017) instead of RGCN (Schlichtkrull et al., 2017) and ignoring edge type during the knowledge graph learning. Without relation learning, we observe a slight decrease in results, emphasizing its importance. Moreover, excluding the GKI component (i.e., without knowledge integration) substantially degrades performance, reinforcing the significance of the GNN mechanism for effective knowledge integration. These experiments highlight the contribution of the complete GKI component to GOProteinGNN’s ability to capture essential biological contexts and interactions, leading to performance improvements.
Method Short Range Contact PPI
P@L P@L/2 P@L/5 F1
GOProteinGNN $\textbf{0.44}^{*}$ $\textbf{0.54}^{*}$ $\textbf{0.66}^{*}$ $\textbf{80.24}^{*}$
-- GKI Ignoring Relations 0.41 0.51 0.64 79.14
-- Without GKI 0.30 0.40 0.52 70.94
Table 5. Ablation test for the GKI component on the contact prediction task, and the PPI task (SHS27K). The metrics are precision scores (P@L, P@L/2, and P@L/5) for short-range contact predictions and the F1 score for the PPI task. The best result is highlighted in bold.
6.2. Number of GKI Components Ablation
We wish to experiment with the number of optimal GKI components. Each GKI component adds an additional knowledge graph injection layer to the model. In the case of GOProteinGNN, the number of GKI components (referred to as $k$) is set to 3. In this ablation test, we reduce the number of GKI components to observe their impact. As shown in Table 6, reducing knowledge injections in GOProteinGNN results in a slight performance decrease. This decrease highlights the significance and meaningfulness of the GKI component.
Method Short Range Contact PPI
P@L P@L/2 P@L/5 F1
GOProteinGNN ($k=3$) $\textbf{0.44}^{*}$ $\textbf{0.54}^{*}$ $\textbf{0.66}^{*}$ $\textbf{80.24}^{*}$
GOProteinGNN ($k=2$) 0.43 0.53 0.64 79.28
Table 6. Ablation test for the number of GKI components ($k$) on the contact prediction task, and the PPI task (SHS27K). The metrics are precision scores (P@L, P@L/2, and P@L/5) for short-range contact predictions and the F1 score for the PPI task. The best result is highlighted in bold.
7. Conclusions
In this paper, we introduced GOProteinGNN, a novel architecture that enhances PLMs by integrating protein knowledge graph information during the creation of amino acid-level representations. Specifically, GOProteinGNN proposed a GKI layer that leverages the [CLS] token representation of the PLM during the encoder stage in a novel manner. Our methodology enables the incorporation of information at both the individual amino acid and protein levels, fostering a comprehensive learning process through graph-based techniques. Furthermore, GOProteinGNN uniquely learns the entire protein knowledge graph during training, allowing it to capture broader relational nuances and dependencies beyond mere triplets as done in previous work. GOProteinGNN showcases its potential across diverse real-world bioinformatics tasks, such as amino acid contact prediction and protein-protein interaction identification. We posit that the observed performance improvements arise from a deeper comprehension of proteins themselves and the associated factual knowledge. This understanding allows generalization for never-before-seen proteins, as demonstrated on downstream tasks. By leveraging both the protein knowledge graph and sequential data modeling, GOProteinGNN effectively propagates information regarding proteins with biological connections. In drug development, it can facilitate virtual screening by prioritizing compounds for binding to specific protein targets, underscoring its potential to provide deeper insights into complex biological phenomena and expediting the drug discovery process.
Appendix A Pre-training Algorithm
1     Input: $\mathcal{G}_{t}$ training set, $\mathcal{G}_{v}$ validation set , $\mathcal{E}_{max}$ maximum epochs, $k$ number of GKI components, $\mathbf{E}$ pretrained encoder (with $L$ layers).
2     Output: $\mathbf{E}$ pretrained model.
3
1:  for $epoch=1$, $2$, ..., $\mathcal{E}_{max}$ do
2:     Sample protein (amino acid sequence) $p\in\mathcal{G}_{t}$
3:     Extract protein knowledge graph $G$
4:     Mask protein: $p=[CLS]\ a_{1}\ a_{2}\ ...\ <MASK>\ ...\ a_{L_{p}}$
5:     $z_{p}=E_{Embedding}(p)\in\mathbb{R}^{L_{p}\times D}$
6:     for $l=1$, $2$, ..., $L$ do
7:        $z_{p}=\mathbf{E}_{l}(z_{p})\in\mathbb{R}^{L_{p}\times D}$
8:        if $l\pmod{L\div k}\equiv 0$ then
9:           $z_{p}[0]=[CLS]^{\prime}_{l}=GKI([CLS]_{l},G)\in\mathbb{R}^{D}$
10:        end if
11:     end for
12:     $\mathcal{L}_{MLM}=-\sum_{j=0}^{M-1}\log P(a_{(j)}|z_{p},\Theta)$
13:     Minimize $\mathcal{L}_{MLM}$ using Adam optimizer
14:     Evaluate the model on $\mathcal{G}_{v}$
15:     if evaluation criterion improves then
16:        Save model
417:     end if
18:  end for
Algorithm 1 GOProteinGNN training algorithm
Appendix B Pre-training Dataset Description
We employed the ProteinKG25 dataset (Zhang et al., 2022) for pre-training, consistent with all knowledge-enhanced baseline methods, such as OntoProtein (Zhang et al., 2022) and KeAP (Zhou et al., 2023). ProteinKG25 provides a knowledge graph with aligned descriptions and protein sequences, respectively, to GO terms (Ashburner et al., 2000) and protein entities. GO terms can be a molecule, a cellular component, or a biological process. The GO terms and their relations to proteins are described using natural language.
The dataset contains around 5 million triplets, including nearly 600k proteins, 50k attribute terms, and 31 relation terms, with an average of 8.64 relations per protein.
Consistent with KeAP (Zhou et al., 2023), we address the data leakage issue in OntoProtein (Zhang et al., 2022) by removing amino acid sequences present in downstream tasks from the knowledge graph, ensuring they are unseen during pre-training. This inductive setup allows for a more accurate assessment of the model’s generalization to novel proteins. KeAP, after mitigating data leakage, is considered SOTA across all tasks. Given the robust performance demonstrated by KeAP, surpassing methods such as OntoProtein, we focus our comparative analysis on KeAP.
B.0.1. Mitigating Limitations in Knowledge Graphs
Knowledge graphs may contain inconsistencies or incomplete data due to their curation process. To address this, we conducted a thorough analysis to detect inconsistencies within GO relations in our dataset, primarily leveraging lexical approaches and synonym exploitation (Mougin, 2015). Our findings demonstrate that the ProteinKG25 dataset is robust, with inconsistencies accounting for less than 0.02% of the triplets. For a fair comparison, these minor inconsistencies were disregarded. Our primary contribution lies in advancing representation learning through a novel knowledge injection approach, building on prior SOTA methods that have shown promising results with existing knowledge graphs. Future work can explore additional strategies to mitigate these limitations further.
Appendix C Experimental Details
C.1. Task Specific Baselines
Building upon the KeAP (Zhou et al., 2023) baselines, we extend our evaluation to encompass additional SOTA models that are known to perform well on specific tasks within the protein domain. For the contact and homology tasks, we include variants of LSTM (Hochreiter and Schmidhuber, 1997), ResNet (He et al., 2015), and Transformer (Vaswani et al., 2017), as well as ProteinBert (Brandes et al., 2021) for the homology. Following Zhang et al., we examine DPPI (Hashemifar et al., 2018), DNNPPI (Li et al., 2018), PIPR (Chen et al., 2019), and GNN-PPI (Lv et al., 2021) for the protein-protein interactions (PPI) task. Also, we incorporate PIPR (Chen et al., 2019) for affinity estimation and MSA Transformer (Rao et al., 2021) and ProtT5-XL (Elnaggar et al., 2022) for semantic similarity inference.
C.2. Implementation Details
During pre-training, GOProteinGNN is trained for 50k steps using a learning rate of 1e-4, weight decay of 0.02, and a batch size of 512 (proteins) over four GPUs (NVIDIA A40, 48GB Memory each). RGCN layers $L$ are set to 1 and GNN Knowledge Injection (GKI) components (referred to as $k$) are set to 3. Our experiments demonstrated that the single-layer RGCN effectively captures most of the critical information, similar to prior studies. We found that adding additional layers did not yield significant gains and, in some cases, introduced noise into the information. For downstream tasks, we follow the SOTA KeAP (Zhou et al., 2023) to ensure a fair comparison. Specifically, we use hyperparameters from GNN-PPI (Lv et al., 2021) for PPI prediction and PROBE (Unsal et al., 2022) for binding affinity and similarity inference. For TAPE (Rao et al., 2019), we performed a grid search to select the optimal hyperparameters. For full implementation details, refer to the provided code.
C.3. GOProteinGNN Computational Complexity
Efficient handling of computational complexity is crucial for graph-based models such as GOProteinGNN, particularly when applied to large-scale protein knowledge graphs.
In terms of computational complexity, GOProteinGNN, like other GCNs, operates at approximately $O(|V|*|E|)$. Several optimizations, including limiting the number of hops per protein, reduce the complexity to $O(|V|)$, achieving performance akin to KeAP. During application, the complexity of both algorithms is $O(1)$.
In our current setup, the full receptive graph (Wang et al., 2023) for each protein is loaded, enabling effective modeling. For scalability with increasing graph size, studies in graph algorithms offer promising solutions, such as dynamically caching receptive graphs (Wang et al., 2023) for individual nodes. These approaches ensure that our method remains practical for larger datasets.
Appendix D Downstream Tasks Definitions
D.0.1. Contact Prediction
Contact prediction is a pairwise amino acid task, where each pair $\langle x_{i},x_{j}\rangle$ of input amino acids from a sequence (protein) $x$ is assigned to a label $y_{ij}\in\{0,1\}$, where the label denotes whether the amino acids are in contact ($<8$Å) (Rao et al., 2019). The data comes from ProteinNet (AlQuraishi, 2019), and the evaluation is based on the precision metric.
D.0.2. Semantic Similarity Inference
This task aims to evaluate how much information representation models capture about biomolecular functional similarity, with a specific emphasis on the biological process (BP) category within the Gene Ontology (GO) dataset. We adopt the semantic similarity measures outlined by Unsal et al. and compute Spearman’s rank correlation between representation vector similarities and the actual GO-based semantic similarities of corresponding protein pairs. Higher correlation values indicate greater success in representing functional similarity.
D.0.3. Protein-Protein Interaction Identification
Protein-protein interactions (PPI) refer to the physical contacts between two or more amino acid sequences, or two proteins, in our study. The goal is to predict the interaction type(s) of each protein pair, out of 7 possible interactions (i.e., multilabel classification task). The evaluation is conducted on the datasets: SHS27K (Chen et al., 2019) and SHS148K (Chen et al., 2019). Breadth-first search (BFS) is used to generate test sets from the two datasets. F1 score is used as the evaluation metric.
D.0.4. Remote Homology Detection
Remote Homology Detection is a sequence-level classification task. Each input sequence (protein) $x$ is assigned a label $y\in\{1,...,1195\}$ representing different possible protein folds. The data is from DeepSF (Hou et al., 2017) and the evaluation is based on average accuracy on a fold-level holdout set.
Appendix E Protein-Protein Binding Affinity Estimation
In this section, we present a task where GOProteinGNN provides limited advantages over existing methods. The Protein-Protein Binding Affinity Estimation task involves evaluating the capability of a protein representation to predict changes in binding affinity resulting from protein mutations. In this context, GOProteinGNN does not incorporate protein structural information and thus offers no additional benefits compared to other methods. The results indicate that while performance is not degraded, it also does not show significant improvement.
E.0.1. Task Definition
Each protein pair is associated with a real value in this task, making it a regression task that reflects binding affinity changes. We follow the approach of Unsal et al. and use the SKEMPI dataset (Moal and Fernández-Recio, 2012) with an evaluation based on the mean squared error obtained by 10-fold cross-validation.
E.0.2. Task Results
Table 7 presents the binding affinity estimation task results. GOProteinGNN outperformed PIPR (Chen et al., 2019) and ProtBert (Elnaggar et al., 2022) with a significant margin of improvement. Furthermore, it is noteworthy that GOProteinGNN exhibited a slight improvement over KeAP (Zhou et al., 2023). The performances of ESM-1b (Rives et al., 2019), ESM-2 (Lin et al., 2023), and GOProteinGNN are comparable, with no statistically significant differences. This similarity in performance may be attributed to the task’s sensitivity to subtle structural changes, which ESM-1b and ESM-2 capture effectively through their extensive pre-training, which are critical for this task (Unsal et al., 2022). In this context, the external factual knowledge leveraged by GOProteinGNN does not provide an advantage. However, the protein knowledge graph includes descriptions of structural properties, allowing GOProteinGNN to inject such knowledge directly into protein representations. This capability contributes to its improved understanding of the protein knowledge graph, leading to modest improvements over KeAP (Zhou et al., 2023).
Method Affinity (↓)
PIPR 0.63
ProtBert 0.58
ESM-1b 0.50
ESM-2 0.50
KeAP 0.52
GOProteinGNN 0.51
Table 7. Results on protein-protein binding affinity prediction task. Affinity (↓) means the lower the better.